OpenStudy (anonymous):

How do you find the integral of sin2x. Is it the sameas sinx?

OpenStudy (anonymous):

\[\frac{-\sin(2x)}{2} +C\]

OpenStudy (anonymous):

\[\frac{-\cos(2x)}{2} +C\] should be this ^ , before i forgot to change the sin to a cos